Calculate the mass of hydrogen formed when 27g of aluminum reacts with excess hydrochloric acod according to the balance equetion belo. 2AL+ 6HCL- 2AlCl3+ 3H2
1
Expert's answer
2021-10-22T03:53:21-0400

2Al + 6HCl → 2AlCl3 + 3H2

M(Al) = 27.0 g/mol

n(Al) "= \frac{27}{27} = 1 \;mol"

According to the reaction:

n(H2) "= \frac{3}{2}n(Al) = \frac{3}{2} \times 1 = 1.5 \;mol"

M(H2) = 2 g/mol

m(H2) "= 1.5 \times 2 = 3 \;g"

Answer: 3 g